# Small-Time Local Controllability of the multi-input bilinear Schrödinger equation thanks to a quadratic term

### v3

@@ A.2.2 Computation of Lie brackets for the Schrödinger equation
If $k=0$, the result in true. Then, assume that $k\geqslant 1$. As $\varphi_{p},\varphi_{q}\in H^{2k}_{(0)}(0,1)$, one has
$\left\langle\underline{\operatorname{ad}}_{A}^{k}(\mu)\varphi_{p},\varphi_{q}\right\rangle=\left\langle[\underline{\operatorname{ad}}_{A}^{k-1}(\mu),A]\varphi_{p},\varphi_{q}\right\rangle=\left\langle(\lambda_{p}I_{d}-A)\underline{\operatorname{ad}}_{A}^{k-1}(\mu)\varphi_{p},\varphi_{q}\right\rangle$
@@
$\left\langle\underline{\operatorname{ad}}_{A}^{k}(\mu)\varphi_{p},\varphi_{q}\right\rangle=\left\langle\underline{\operatorname{ad}}_{A}^{k-1}(\mu)\varphi_{p},(\lambda_{p}I_{d}-A)\varphi_{q}\right\rangle=(\lambda_{p}-\lambda_{q})\left\langle\underline{\operatorname{ad}}_{A}^{k-1}(\mu)\varphi_{p},\varphi_{q}\right\rangle.$
An induction gives the result. ∎
@@ A.3 Some elements of proof in finite dimension
In this section, we present a proof of Theorem 2.15 relying on Sussmann’s $S(\theta)$-sufficient condition (see [45, Theorem 7.3]), recalled in Proposition A.12 below (see [28, Theorem 3.29]), for which we need the following definition.
@@
then the system $\dot{x}=f_{0}(x)+uf_{1}(x)+vf_{2}(x)$ is $L^{\infty}$-STLC.
@@
Initialization for $m=0$: we prove $L^{\infty}$-STLC by applying Proposition A.12. First (2.7) gives the Lie algebra rank condition,
$\text{Span}(f_{b_{1}}(0),\cdots f_{b_{d}}(0))=\mathbb{R}^{d},$
with $b_{1},\cdots,b_{r}\in\mathcal{B}_{1}$ such that $\text{Span}(f_{b_{1}}(0),\cdots f_{b_{r}}(0))=S_{1}(f)(0)$ and $b_{r+1},\cdots,b_{d}\in\mathcal{B}_{2,good}$. One considers
@@
    Either, $n(\mathfrak{b})\geqslant 4$ and then (2.7) gives the compensation because, for all $i\in\llbracket 1,d\rrbracket$,
    $n(b_{i})+\theta n_{0}(b_{i})\leqslant 2+\theta\left(\underset{i\in\llbracket 1,d\rrbracket}{\max}|b_{i}|-1\right)<3<4+\theta n_{0}(\mathfrak{b})\leqslant n(\mathfrak{b})+\theta n_{0}(\mathfrak{b}).$
@@
    Or, $n(\mathfrak{b})=2$ i.e. $(n_{1}(\mathfrak{b}),n_{2}(\mathfrak{b}))\in\{(2,0),(0,2)\}$ i.e. $\sigma(\mathfrak{b})\in\text{Span}\{\mathcal{B}_{2,bad}\}$. Thus:
@@
      $n(b_{i})+\theta n_{0}(b_{i})=1+\theta n_{0}(b_{i})<2\leqslant n(\mathfrak{b})+\theta n_{0}(\mathfrak{b}).$
@@
    Theses points prove (A.9).
Heredity. We assume the result is proved up to some $m\in\mathbb{N}$. We consider the extended system
$\left\{\begin{array}[]{l}\dot{x}=f_{0}(x)+uf_{1}(x)+vf_{2}(x)\\ \dot{u}=w\\ \dot{v}=\omega\end{array}\right.$
with state $Y=(x,u,v)\in\mathbb{R}^{d+2}$ and control $(w,\omega)$, i.e. $\frac{\mathrm{d}Y}{\mathrm{d}t}=F_{0}(Y)+wF_{1}(Y)+\omega F_{2}(Y)$ with:
@@
$\forall j\geqslant 1\,,\quad F_{M_{j}^{i}}(Y)=\begin{pmatrix}\underline{\operatorname{ad}}_{f_{0}+uf_{1}+vf_{2}}^{j-1}(f_{i})(x)\\ 0\\ 0\end{pmatrix},$
where $\underline{\operatorname{ad}}$ denotes the right-bracketing: $\underline{\operatorname{ad}}_{f}(g)=[g,f]$. Thus, for every $i\in\llbracket 1,2\rrbracket$ and $l\in\mathbb{N}$, $F_{W_{1,l}^{i}}=0$ and for every $j\geqslant 2$,
@@
Moreover, for every $l\in\mathbb{N}$, $F_{C_{0,l}}=F_{C_{1,l}}=0$ and, for every $j\geqslant 2$,
$F_{C_{j,l}}(Y)=\begin{pmatrix}(-1)^{j}\underline{\operatorname{ad}}_{f_{0}+uf_{1}+vf_{2}}^{l}[\underline{\operatorname{ad}}_{f_{0}+uf_{1}+vf_{2}}^{\left\lfloor\frac{j+1}{2}\right\rfloor-1}(f_{1}),\underline{\operatorname{ad}}_{f_{0}+uf_{1}+vf_{2}}^{\left\lfloor\frac{j}{2}\right\rfloor-1}(f_{2})](x)\\ 0\\ 0\end{pmatrix}.$
@@
Moreover, for every $i\in\llbracket 1,2\rrbracket$, $j\geqslant 2$, $l\in\mathbb{N}$,
$F_{W_{j,l}^{i}}(0)=\begin{pmatrix}f_{W_{j-1,l}^{i}}(0)\\ 0\\ 0\end{pmatrix},\qquad F_{C_{j,l}}(0)=\begin{pmatrix}f_{C_{j-1,l}}(0)\\ 0\\ 0\end{pmatrix}.$
Thus the extended system satisfies the assumptions of Theorem 2.15 with $d+2\to d$. By the induction assumption, the extended system is $W^{m,\infty}$-STLC. By choosing trajectories of the extended system starting and finishing at $u=0$, one concludes that the initial system $\dot{x}=f_{0}(x)+uf_{1}(x)+vf_{2}(x)$ is $W^{m+1,\infty}$-STLC. ∎
@@
If $\mathcal{B}\subset\operatorname{Br}(X)$ is a Hall set, there exists a unique family $(\eta_{b})_{b\in\mathcal{B}}$ of maps +×$\mathbb{R}_{+}\times L^{1}_{loc}(\mathbb{R}_{+},\mathbb{R})^{2}\rightarrow\mathbb{R}$, called coordinates of the pseudo first kind associated with $\mathcal{B}$ and they satisfy the announced homogeneity properties in (2.16). Moreover, by [9, Section 2.5], for every $t>0$ and $u,v\in L^{1}((0,t),\mathbb{R})$,
@@ A.4 Some lemmas about ODEs
Let $\delta>0$ and $z\in\mathcal{C}^{1}(B(0,{\delta});\mathbb{R}^{d})$ be such that $\left\|z\right\|_{\infty}\leqslant\delta$. We note $x(\cdot,z,0)$ the solution to the system: $\left\{\begin{matrix}x^{\prime}&=&z(x)\\ x(0)&=&0\end{matrix}\right.$. Then,
@@
This result is linked to continuity with respect to the initial condition ∎
@@
Under the same assumptions as the previous lemma, there exists $r(\delta)>0$, $C(\delta)>0$, such that, for all $T_{1}<T<T_{1}+1$, $p\in B\left(0,r(\delta)\right)$, and $u,v\in L^{\infty}(T_{1},T)$ with $\left\|u\right\|_{L^{\infty}},\left\|v\right\|_{L^{\infty}}<r(\delta)$, one has
@@
This statement is based on the previous Lemma and the Grönwall inequality. ∎
@@
Let $p\in B(0,\varepsilon)$ with $\varepsilon=\min\left(r(\delta),r(h)\right)$. Using the Duhamel formula,
@@
with $z(T_{1})=0$. We estimate each term of this formulation
@@
The Grönwall lemma gives the result. ∎
@@ Acknowledgement
The author acknowledges support from grants ANR-20-CE40-0009fh (Project TRECOS) and ANR-11-LABX-0020 (Labex Lebesgue), as well as from the Fondation Simone et Cino Del Duca – Institut de France.

### v4

@@ A.2.2 Computation of Lie brackets for the Schrödinger equation
Let $k\in\mathbb{N}$. As $\varphi_{p},\varphi_{q}\in H^{2k+2}_{(0)}(0,1)$, one has
$\left\langle\underline{\operatorname{ad}}_{A}^{k+1}(\mu)\varphi_{p},\varphi_{q}\right\rangle=\left\langle[\underline{\operatorname{ad}}_{A}^{k}(\mu),A]\varphi_{p},\varphi_{q}\right\rangle=\left\langle(\lambda_{p}I_{d}-A)\underline{\operatorname{ad}}_{A}^{k}(\mu)\varphi_{p},\varphi_{q}\right\rangle.$
@@
$\left\langle\underline{\operatorname{ad}}_{A}^{k+1}(\mu)\varphi_{p},\varphi_{q}\right\rangle=\left\langle\underline{\operatorname{ad}}_{A}^{k}(\mu)\varphi_{p},(\lambda_{p}I_{d}-A)\varphi_{q}\right\rangle=(\lambda_{p}-\lambda_{q})\left\langle\underline{\operatorname{ad}}_{A}^{k}(\mu)\varphi_{p},\varphi_{q}\right\rangle.$
Moreover, the result in true for $k=0$. An induction concludes the proof. ∎
@@ A.3 Some elements of proof in finite dimension
In this section, we present a proof of Theorem 2.15 relying on Sussmann’s $S(\theta)$-sufficient condition (see [46, Theorem 7.3]), recalled in Proposition A.13 below (see [28, Theorem 3.29]), for which we need the following definition.
@@
then the system $x^{\prime}=f_{0}(x)+uf_{1}(x)+vf_{2}(x)$ is $L^{\infty}$-STLC.
@@
Initialization $(m=0)$: we prove $L^{\infty}$-STLC by applying Proposition A.13. First (2.7) gives the Lie algebra rank condition $\text{Span}(f_{b_{1}}(0),\cdots f_{b_{d}}(0))=\mathbb{R}^{d},$ with $b_{1},\cdots,b_{r}\in\mathcal{B}_{1}$ such that $\text{Span}(f_{b_{1}}(0),\cdots f_{b_{r}}(0))=S_{1}(f)(0)$ and $b_{r+1},\cdots,b_{d}\in\mathcal{B}_{2,good}$. One considers
@@
    either $n(\mathfrak{b})\geqslant 4$ and then (2.7) gives the compensation because, for all $i\in\llbracket 1,d\rrbracket$,
    $n(b_{i})+\theta n_{0}(b_{i})\leqslant 2+\theta\left(\underset{i\in\llbracket 1,d\rrbracket}{\max}|b_{i}|-1\right)<3<4+\theta n_{0}(\mathfrak{b})\leqslant n(\mathfrak{b})+\theta n_{0}(\mathfrak{b}),$
@@
    or $n(\mathfrak{b})=2$ i.e. $(n_{1}(\mathfrak{b}),n_{2}(\mathfrak{b}))\in\{(2,0),(0,2)\}$ i.e. $\sigma(\mathfrak{b})\in\text{Span}\{\mathcal{B}_{2,bad}\}$. Thus
@@
      $n(b_{i})+\theta n_{0}(b_{i})=1+\theta n_{0}(b_{i})<2\leqslant n(\mathfrak{b})+\theta n_{0}(\mathfrak{b}),$
@@
Theses points prove (A.8). The system $x^{\prime}=f_{0}(x)+uf_{1}(x)+vf_{2}(x)$ is $L^{\infty}-$STLC.
Heredity: we assume that the result is proved up to some $m\in\mathbb{N}$. We consider the extended system
$\left\{\begin{array}[]{l}\dot{x}=f_{0}(x)+uf_{1}(x)+vf_{2}(x)\\ \dot{u}=w\\ \dot{v}=\omega\end{array}\right.,$
with state $Y=(x,u,v)\in\mathbb{R}^{d+2}$ and control $(w,\omega)$, i.e. $Y^{\prime}=F_{0}(Y)+wF_{1}(Y)+\omega F_{2}(Y)$ with
@@
$\forall i\in\llbracket 1,2\rrbracket,\ \forall j\geqslant 1,\quad F_{M_{j}^{i}}(Y)=\begin{pmatrix}\underline{\operatorname{ad}}_{f_{0}+uf_{1}+vf_{2}}^{j-1}(f_{i})(x)\\ 0\\ 0\end{pmatrix}.$
Thus, for every $i\in\llbracket 1,2\rrbracket$ and $l\in\mathbb{N}$, $F_{W_{1,l}^{i}}=0$ and for every $i\in\llbracket 1,2\rrbracket$, $j\geqslant 2$, $l\in\mathbb{N}$,
@@
Moreover, for every $l\in\mathbb{N}$, $F_{C_{0,l}}=F_{C_{1,l}}=0$ and, for every $j\geqslant 2$, $l\in\mathbb{N}$,
$F_{C_{j,l}}(Y)=\begin{pmatrix}(-1)^{j}\underline{\operatorname{ad}}_{f_{0}+uf_{1}+vf_{2}}^{l}\left[\underline{\operatorname{ad}}_{f_{0}+uf_{1}+vf_{2}}^{\left\lfloor\frac{j+1}{2}\right\rfloor-1}(f_{1}),\underline{\operatorname{ad}}_{f_{0}+uf_{1}+vf_{2}}^{\left\lfloor\frac{j}{2}\right\rfloor-1}(f_{2})\right](x)\\ 0\\ 0\end{pmatrix}.$
@@
Moreover, for every $j\geqslant 2$, $l\in\mathbb{N}$,
$F_{W_{j,l}^{i}}(0)=\begin{pmatrix}f_{W_{j-1,l}^{i}}(0)\\ 0\\ 0\end{pmatrix},\qquad F_{C_{j,l}}(0)=\begin{pmatrix}f_{C_{j-2,l}}(0)\\ 0\\ 0\end{pmatrix}.$
Thus, the extended system satisfies the assumptions of Theorem 2.15 with dimension $d+2$. By the induction hypothesis, the extended system is $W^{m,\infty}$-STLC. By choosing trajectories of the extended system starting and finishing at $u,v=0$, one concludes that the initial system $x^{\prime}=f_{0}(x)+uf_{1}(x)+vf_{2}(x)$ is $W^{m+1,\infty}$-STLC. ∎
@@
If $\mathcal{B}\subset\operatorname{Br}(X)$ is a Hall set, there exists a unique family $(\eta_{b})_{b\in\mathcal{B}}$ of maps +×$\mathbb{R}_{+}\times L^{1}_{loc}(\mathbb{R}_{+},\mathbb{R})^{2}\rightarrow\mathbb{R}$, called coordinates of the pseudo-first kind associated with $\mathcal{B}$ and they satisfy the announced homogeneity properties in (2.16). Moreover, by [9, Section 2.5], for every $t>0$ and $u,v\in L^{1}((0,t),\mathbb{R})$,
@@ A.4 Some lemmas about ODEs
Let $\delta>0$ and $z\in\mathcal{C}^{1}(B(0,{\delta}),\mathbb{R}^{d})$ be such that $\left\|z\right\|_{\infty}\leqslant\delta$. We note $x(\cdot;z,0)$ the solution to the system: $\left\{\begin{matrix}x^{\prime}&=&z(x)\\ x(0)&=&0\end{matrix}\right.$. Then,
@@
This result is linked to continuity with respect to the initial condition. ∎
@@
Let $\delta>0$, and $f_{0},f_{1},f_{2}:B(0,2\delta)\to\mathbb{R}^{d}$ be real continuous functions, with $f_{0}(0)=0$. There exists $r(\delta),C(\delta)>0$, such that, for all $T_{1}<T<T_{1}+1$, $p\in B\left(0,r(\delta)\right)$, and $u,v\in L^{\infty}(T_{1},T)$ with $\left\|u\right\|_{L^{\infty}},\left\|v\right\|_{L^{\infty}}<r(\delta)$, one has
@@
This statement is based on the previous lemma and the Grönwall’s inequality. ∎
@@
Let $p\in B(0,\varepsilon)$ with $\varepsilon=\min\left(r(\delta),r(h)\right)$. Using the Duhamel’s principle,
@@
with $z(T_{1})=0$. We estimate each term of this decomposition:
@@
The Grönwall’s lemma gives the result. ∎
@@ Acknowledgement
The author acknowledges support from grants ANR-20-CE40-0009 (Project TRECOS) and ANR-11-LABX-0020 (Labex Lebesgue), as well as from the Fondation Simone et Cino Del Duca – Institut de France.